the local theater has three types of seats for broadway plays: main floor, balcony, and mezzanine. Main floor tickets are $54, balcony are $37, and mezzanine are $27. one particular night sales totaled $58,921. There were 350 more main floor tickets sold than balcony and mezzanine sold combined. the number of balcony tickets sold is 43 more than the 3 times the number of mezzanine how many of each types of tickets were sold
Answer by Boreal(15235)

Mezzanine=x
Balcony=3x+43
Main floor=4x+43+350, or 4x+393
54(4x+393)+37(3x+43)+27x=58921
216x+21222+111x+1591+27x=58921
354x+22813=58921
354x=36108
x=102 Mezzanine or $2754
3x+43=Balcony =349 for $12913
Main floor=4x+393 = 801 for $43254